# Analytics, have some humility: a statistical view of fourth-down decision making

### v5

@@ 1 Introduction
The two most widely used value functions by analysts of American football are win probability and expected points (Yurko et al.,, 2018). Win probability (${\mathsf{WP}}$) measures the probability that the team with possession at the current game-state wins the game. Expected points (${\mathsf{EP}}$) measures the expected value of the net number of points of the next score in the game, relative to the team with possession, given the current game-state. The most prominent example of analysts using these value functions to dictate in-game strategy is fourth-down decision making. On fourth down, a football coach has three choices: go for it (${\mathsf{Go}}$), attempt a field goal (${\mathsf{FG}}$), or punt the ball (${\mathsf{Punt}}$). Previous work by Romer, (2006) and Burke, (2009) suggest making the decision that maximizes estimated ${\mathsf{EP}}$. Yet a team’s goal is to win the game, not score more points on average, so ${\mathsf{EP}}$ is the wrong objective function. Hence, modern approaches by Baldwin, (2021) and Burke111 Burke releases fourth-down decision recommendations with the ESPN analytics team, and their recommendation algorithm is proprietary. suggest making the decision that maximizes estimated ${\mathsf{WP}}$. Each of analyses found that National Football League (NFL) coaches are too conservative on fourth down; they often settle for kicks even when they should go for it.
@@ 2.1 Data
We access every NFL play from 1999 to 2022 using the ${\mathsf{R}}$ package nflFastR (Carl and Baldwin,, 2022). Each play includes variables that describe the context of the play, which are relevant to estimating win probability, such as the score differential, time remaining, yardline, down, yards to go, etc. (see Table 2 in Appendix A for descriptions of relevant variables). Note that yardline is a real number in $\{0,1,...,99,100\}$ measuring the number of yards from the opponent’s endzone, where 00 represents a touchdown and $100$ indicates a safety. Holding out plays from 1999 to 2005 and from 2022 for various models and validation procedures discussed later in this paper, we are left with a primary dataset of $600,825$ plays, with $229,635$ first-down plays and $4,101$ non-tied games from 2006 to 2021, henceforth referred to as the “observed play-by-play football dataset.” We fit our win probability models using this dataset. The code for this study, which includes code to scrape the dataset used in this study, is publicly available on Github.222 https://github.com/snoopryan123/fourth_down
@@ 2.2 Example plays
Figure 2: Burke’s decision boundary chart for example play 2. The chart visualizes the estimated optimal decision according to effect size (color) as a function of yardline ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. The yellow dot denotes the actual play’s yardline and yards to go.
@@ 2.3 Estimating the win probability of a fourth-down decision
We model the probability of a successful field goal as a function of yardline and kicker quality in Section 2.5. If the kicking team misses the field goal and the spot of the kick is within the 20 yardline, the opposing team has a first-down possession at yardline $80$. If the kicking team misses the field goal and the spot of the kick is beyond the 20 yardline, the opposing team has a first-down possession at the spot of the kick. The spot of the kick is typically 7 yards behind the line of scrimmage, so in terms of the current yardline $y$, the next yardline relative to the opposing team is $\min\{80,100-(y+7)\}$. Denoting the remaining game-state variables relative to the opposing team at the next play by ${\mathbf{x}}^{\prime}$, we have
@@
The vast majority of kickoffs end in a touchback (yardline $75$), so for simplicity we instead compute
@@ 2.4 Estimating first-down win probability
Lock and Nettleton, (2014) use a Random Forest (${\mathsf{RF}}$) (Breiman,, 2001) to estimate win probability from historical data. The response variable is a binary variable indicating whether the team with possession wins the game. They model win probability as a function of score differential, game seconds remaining, yardline, down, yards to go, the number of timeouts remaining for each team, pre-game point spread, total points scored, and an additional feature to capture the change in impact of score differential over the course of a game,
@@
Baldwin, (2021) uses ${\mathsf{XGBoost}}$ (Chen and Guestrin,, 2016) to estimate win probability from historical data. He uses the same binary win/loss response variable as before. Baldwin models win probability as a function of score differential, game seconds remaining, half seconds remaining, yardline, down, yards to go, whether the team with possession is at home, whether the team with possession receives the second half kickoff, and the number of timeouts remaining for each team. He uses two additional features to capture the change in impact of point spread and score differential over the course of a game,
@@
Baldwin includes monotonic constraints for yardline, yards to go, down, score differential, timeouts remaining for each team, spread-time, and diff-time-ratio. He tunes the ${\mathsf{XGBoost}}$ hyper-parameters by minimizing cross validated log-loss (Baldwin,, 2021).
@@
The aforementioned models are fit from a historical play-by-play dataset that includes plays from all downs. Those models fit first-down win probability by including down as a covariate. We fit an ${\mathsf{XGBoost}}$ model just using first-down plays. This model predicts binary win/loss as a function of score differential, game seconds remaining, pre-game point spread, yardline, receive $2^{nd}$ half kickoff indicator, offensive team’s number of timeouts remaining, defensive team’s number of timeouts remaining, total score, and
@@
We include monotonic increasing constraints for score differential, scoreTimeRatio, and offensive timeouts remaining and monotonic decreasing constraints for point spread, yardline, and defensive timeouts remaining.
Now, we compare the out-of-sample predictive performance of various first-down win probability models. Our full dataset consists of all football plays from 2006 to 2021. We split our dataset in half by randomly sampling 50% of the games. The first-down plays from the first 50% of these games form the hold-out test set. All the plays from the other 50% of these games form the training set. To tune the ${\mathsf{XGBoost}}$ models, we split the training set in half by randomly sampling 50% of the games from the training set. The plays from the first 50% of these games form the ${\mathsf{XGBoost}}$ training set, and the remaining plays form the validation set for hyper-parameter tuning. We then tune our ${\mathsf{XGBoost}}$ models in a similar fashion as Baldwin, (2021). We view the results of our prediction contest in Table 1. Our model fit from just first-down plays performs the best out-of-sample. Thus, in the remainder of this paper we use our proposed first-down win probability model.
model name model type plays in dataset out-of-sample ${\mathsf{logloss}}$
Our proposed model ${\mathsf{XGBoost}}$ first downs 0.440
Lock and Nettleton, (2014) ${\mathsf{RF}}$ all downs 0.447
Baldwin, (2021) ${\mathsf{XGBoost}}$ all downs 0.476
@@ 2.5 Estimating fourth-down decision transition probabilities
Fitting this model on our dataset of $15,472$ observed field goals from 2006 to 2021 yields nontrivial probability predictions for extremely long field goals that have never before been made (e.g., nontrivial probability for a 73-yard field goal from the 55 yardline). To shrink these field goal probability predictions to zero, we impute 500 synthetic missed field goals, randomly distributed from the 51 to the 99 yardline, into our dataset. We visualize this model in Figure 3b.
@@
We use logistic regression to model fourth-down conversion probability as a function of yards to go, down (third vs. fourth down), and our offensive and defensive quality metrics (whose specification is detailed in Appendix B): quarterback quality of the offensive team (${\mathsf{qbqot}}$), non-quarterback offensive quality of the offensive team (${\mathsf{oqrot}}$), defensive quality of the defensive team against the pass (${\mathsf{dqdtap}}$), and defensive quality of the defensive team against the run (${\mathsf{dqdtar}}$). We use these team quality metrics, rather than pre-game point spread, since they yield better predictive performance. This may be because quarterback quality significantly impacts conversion probability while the other team quality measures have a small impact, as seen in Figure 4. Formally, we model
$\begin{split}\log\bigg{(}\frac{{\mathbb{P}}({\text{convert}})}{1-{\mathbb{P}}({\text{convert}})}\bigg{)}=&\ \vec{\alpha_{1}}\cdot\mathbb{I}\left(\text{fourth down}\right)\cdot\text{B-spline}\big{(}\log(\text{yards to go}+1),\ \text{df}=4\big{)}\\ +&\ \vec{\alpha_{2}}\cdot\mathbb{I}\left(\text{third down}\right)\cdot\text{B-spline}\big{(}\log(\text{yards to go}+1),\ \text{df}=4\big{)}\\ +&\ \beta_{0}+\beta_{1}\cdot{\mathsf{qbqot}}+\beta_{2}\cdot{\mathsf{oqrot}}+\beta_{3}\cdot{\mathsf{dqdtap}}+\beta_{4}\cdot{\mathsf{dqdtar}}.\end{split}$ (17)
Figure 4: Fourth-down conversion probability ($y$-axis) according to our model as a function of yards to go ($x$-axis) and team quality (color). Quarterback quality has by far the largest impact on conversion probability, visualized by a larger spread amongst the colored lines.
@@
$\begin{split}{\mathbb{E}}_{\text{go}}[\Delta|{\mathbf{x}},\Delta\geq z]=&\ \vec{\alpha_{1}}\cdot\mathbb{I}\left(\text{fourth down}\right)\cdot\text{B-spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \vec{\alpha_{2}}\cdot\mathbb{I}\left(\text{third down}\right)\cdot\text{B-spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \vec{\alpha_{3}}\cdot\mathbb{I}\left(\text{yards to go}=1\right)\cdot\text{B-spline}\big{(}\text{yardline},\ \text{df}=3\big{)}\\ +&\ \vec{\alpha_{4}}\cdot\mathbb{I}\left(\text{yards to go}\neq 1\right)\cdot\text{B-spline}\big{(}\text{yardline},\ \text{df}=4\big{)}\\ +&\ \beta_{0}+\beta_{1}\cdot{\mathsf{qbqot}}+\beta_{2}\cdot{\mathsf{oqrot}}+\beta_{3}\cdot{\mathsf{dqdtap}}+\beta_{4}\cdot{\mathsf{dqdtar}}.\end{split}$ (18)
@@
$\begin{split}{\mathbb{E}}_{\text{go}}[\Delta|{\mathbf{x}},\Delta<z]=&\ \alpha_{1}\cdot\mathbb{I}\left(\text{fourth down}\right)\cdot\log(\text{yards to go}+1)\\ +&\ \alpha_{2}\cdot\mathbb{I}\left(\text{third down}\right)\cdot\log(\text{yards to go}+1)\\ +&\ \beta_{0}+\beta_{1}\cdot{\mathsf{qbqot}}+\beta_{2}\cdot{\mathsf{oqrot}}+\beta_{3}\cdot{\mathsf{dqdtap}}+\beta_{4}\cdot{\mathsf{dqdtar}}.\end{split}$ (19)
@@ 3 Fourth-down decision making under uncertainty
Moreover, the binary win/loss response variable is clustered: all plays from the same game share the same win/loss outcome. This dependence structure reduces the effective sample size, which is somewhere between the number of first-down plays ($229,635$) and the number of non-tied games ($4,101$) in our dataset. According to the simulation study in Brill and Wyner, (2024), we effectively have half as much data as we think. In other words, a win probability estimator fit from a dataset consisting of half as many plays, but where each play has an independent outcome, has the same accuracy as an estimator fit from a clustered dataset like our historical play-by-play dataset.
Thus, there is likely not enough data to experience the full variability of the nonlinear and interacting variables of score differential, time remaining, point spread, yardline, yards to go, timeouts, etc. In fitting win probability models, we are in a limited-data context. Therefore, it is imperative to incorporate uncertainty quantification into the fourth-down decision procedure.
Since we estimate win probability using machine learning models, we need a non-parametric method to quantify uncertainty in these estimates. Bootstrapping is a natural choice to capture such uncertainty. The bootstrapping process begins with generating $B$ bootstrapped datasets from our original dataset. To do so, we use a randomized cluster bootstrap: re-sample games uniformly with replacement and within each game re-sample drives uniformly with replacement. According to the simulation study in Brill and Wyner, (2024), although the randomized cluster bootstrap achieves higher coverage than the standard i.i.d. bootstrap (re-sample plays uniformly with replacement), it produces undercovered intervals that are too narrow. Thus, we also tune our randomized cluster bootstrap using a fractional bootstrap as recommended in Brill and Wyner, (2024). We choose an appropriate value for $B$ ($B=101$) in Appendix C.
@@
Regardless of the effect size, when ${\mathsf{boot\%}}$ is low we cannot trust the point estimate: the effect size at that game-state is too dependent on the random idiosyncrasies of its particular training set. In other words, the edge detected by the point estimate is more due to noise than signal. In a “parallel universe” in which we re-drew each random win/loss game outcome, the estimated optimal decision at that game-state may differ solely due to the randomness inherent in drawing the history of football.

### v6

@@ 1 Introduction
The two most widely used value functions by analysts of American football are win probability and expected points (Yurko et al.,, 2018). Win probability (${\mathsf{WP}}$) measures the probability that the team with possession at the current game-state wins the game. Expected points (${\mathsf{EP}}$) measures the expected value of the net number of points of the next score in the game, relative to the team with possession, given the current game-state. The most prominent example of analysts using these value functions to dictate in-game strategy is fourth-down decision making. On fourth down, a football coach has three choices: go for it (${\mathsf{Go}}$), attempt a field goal (${\mathsf{FG}}$), or punt the ball (${\mathsf{Punt}}$). Previous work by Romer, (2006) and Burke, (2009) suggest making the decision that maximizes estimated ${\mathsf{EP}}$. Yet a team’s goal is to win the game, not score more points on average, so ${\mathsf{EP}}$ is the wrong objective function. Hence, modern approaches by Baldwin, (2021) and Burke111 Burke releases fourth-down decision recommendations with the ESPN analytics team, and their recommendation algorithm is proprietary. suggest making the decision that maximizes estimated ${\mathsf{WP}}$. Each of these analyses found that National Football League (NFL) coaches are too conservative on fourth down; they often settle for kicks even when they should go for it.
@@ 2.1 Data
We access every NFL play from 1999 to 2022 using the ${\mathsf{R}}$ package nflFastR (Carl and Baldwin,, 2022). Each play includes variables that describe the context of the play, which are relevant to estimating win probability, such as the score differential, time remaining, yards to opponent endzone (i.e., yardline), down, yards to go, etc. (see Table 2 in Appendix A for descriptions of relevant variables). Note that yards to opponent endzone is an integer in $\{0,1,...,99,100\}$, where 00 represents a touchdown and $100$ indicates a safety. Holding out plays from 1999 to 2005 and from 2022 for various models and validation procedures discussed later in this paper, we are left with a primary dataset of $600,825$ plays, with $229,635$ first-down plays and $4,101$ non-tied games from 2006 to 2021, henceforth referred to as the “observed play-by-play football dataset.” We fit our win probability models using this dataset. The code for this study, which includes code to scrape the dataset used in this study, is publicly available on Github.222 https://github.com/snoopryan123/fourth_down
@@ 2.2 Example plays
Figure 2: Burke’s decision boundary chart for example play 2. The chart visualizes the estimated optimal decision according to effect size (color) as a function of yards to opponent endzone ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. The yellow dot denotes the actual play’s yards to opponent endzone and yards to go.
@@ 2.3 Estimating the win probability of a fourth-down decision
We model the probability of a successful field goal as a function of yardline and kicker quality in Section 2.5. If the kicking team misses the field goal and the spot of the kick is within 20 yards to the opponent’s endzone, the opposing team has a first-down possession at $80$ yards to the opponent’s endzone. If the kicking team misses the field goal and the spot of the kick is beyond 20 yards to the opponent’s endzone, the opposing team has a first-down possession at the spot of the kick. The spot of the kick is typically 7 yards behind the line of scrimmage, so in terms of the current yardline $y$, the next yardline relative to the opposing team is $\min\{80,100-(y+7)\}$. Denoting the remaining game-state variables relative to the opposing team at the next play by ${\mathbf{x}}^{\prime}$, we have
@@
The vast majority of kickoffs end in a touchback (yardline $75$),555 It is worth noting that the new NFL kickoff rule will change this (see https://operations.nfl.com/the-rules/rules-changes/dynamic-kickoff-rule-explainer/). so for simplicity we instead compute
@@ 2.4 Estimating first-down win probability
Lock and Nettleton, (2014) use a Random Forest (${\mathsf{RF}}$) (Breiman,, 2001) to estimate win probability from historical data. The response variable is a binary variable indicating whether the team with possession wins the game. They model win probability as a function of score differential, game seconds remaining, yards to opponent endzone, down, yards to go, the number of timeouts remaining for each team, pre-game point spread, total points scored, and an additional feature to capture the change in impact of score differential over the course of a game,
@@
Baldwin, (2021) uses ${\mathsf{XGBoost}}$ (Chen and Guestrin,, 2016) to estimate win probability from historical data. He uses the same binary win/loss response variable as before. Baldwin models win probability as a function of score differential, game seconds remaining, half seconds remaining, yards to opponent endzone, down, yards to go, whether the team with possession is at home, whether the team with possession receives the second half kickoff, and the number of timeouts remaining for each team. He uses two additional features to capture the change in impact of point spread and score differential over the course of a game,
@@
Baldwin includes monotonic constraints for yards to opponent endzone, yards to go, down, score differential, timeouts remaining for each team, spread-time, and diff-time-ratio. He tunes the ${\mathsf{XGBoost}}$ hyper-parameters by minimizing cross validated log-loss (Baldwin,, 2021).
@@
The aforementioned models are fit from a historical play-by-play dataset that includes plays from all downs. Those models fit first-down win probability by including down as a covariate. We fit an ${\mathsf{XGBoost}}$ model just using first-down plays. This model predicts binary win/loss as a function of score differential, game seconds remaining, pre-game point spread, yards to opponent endzone, receive $2^{nd}$ half kickoff indicator, offensive team’s number of timeouts remaining, defensive team’s number of timeouts remaining, total score, and
@@
We include monotonic increasing constraints for score differential, scoreTimeRatio, and offensive timeouts remaining and monotonic decreasing constraints for point spread, yards to opponent endzone, and defensive timeouts remaining.
Now, we compare the out-of-sample predictive performance of various first-down win probability models. Our full dataset consists of all football plays from 2006 to 2021. We split our dataset in half by randomly sampling 50% of the games. The first-down plays from the first 50% of these games form the hold-out test set. All the plays from the other 50% of these games form the training set. To tune the ${\mathsf{XGBoost}}$ models, we split the training set in half by randomly sampling 50% of the games from the training set. The plays from the first 50% of these games form the ${\mathsf{XGBoost}}$ training set, and the remaining plays form the validation set for hyper-parameter tuning. We then tune our ${\mathsf{XGBoost}}$ models in a similar fashion as Baldwin, (2021). We view the results of our prediction contest in Table 1. For reference, we include the predictive performance of two baseline models, a fair coin (always predict 0.5) and a game-level logistic regression model with one linear term for pre-game point spread. Each model’s reduction in error (relative to a fair coin) is the negative percent difference between it’s accuracy and the fair coin’s. Our model fit from just first-down plays performs the best out-of-sample. Thus, in the remainder of this paper we use our proposed first-down win probability model.
model name model type plays in training data out-of-sample ${\mathsf{logloss}}$ reduction in error
Our proposed model ${\mathsf{XGBoost}}$ first downs 0.440 36.54%
Lock and Nettleton, (2014) ${\mathsf{RF}}$ all downs 0.446 36.50%
Baldwin, (2021) ${\mathsf{XGBoost}}$ all downs 0.476 31.27%
Pre-game point spread only logistic regression 0.609 12.10%
Fair coin a constant 0.693 0%
@@ 2.5 Estimating fourth-down decision transition probabilities
Fitting this model on our dataset of $15,472$ observed field goals from 2006 to 2021 yields nontrivial probability predictions for extremely long field goals that have never before been made (e.g., nontrivial probability for a 73-yard field goal from the 55 yards to the opponent’s endzone). To shrink these field goal probability predictions to zero, we impute 500 synthetic missed field goals, randomly distributed from the 51 to 99 yards to the opponent’s endzone, into our dataset. We visualize this model in Figure 3b.
@@
We use logistic regression to model fourth-down conversion probability as a function of yards to go, down (third vs. fourth down), and a market-derived measure of how much better the offensive team’s offensive quality is than the defensive team’s defensive quality (denoted $\Delta{\mathsf{TQ}}$, or the difference in team quality, whose specification is detailed in Appendix B). Formally, we model
$\log\bigg{(}\frac{{\mathbb{P}}({\text{convert}})}{1-{\mathbb{P}}({\text{convert}})}\bigg{)}=$ $\ \vec{\alpha_{1}}\cdot\mathbb{I}\left(\text{fourth down}\right)\cdot\text{B-spline}\big{(}\log(\text{yards to go}+1),\ \text{df}=4\big{)}$ (17)
$+$ $\ \vec{\alpha_{2}}\cdot\mathbb{I}\left(\text{third down}\right)\cdot\text{B-spline}\big{(}\log(\text{yards to go}+1),\ \text{df}=4\big{)}$ (18)
$+$ $\ \beta_{0}+\beta_{1}\cdot\Delta{\mathsf{TQ}}.$ (19)
We visualize this model in Figure 4.
Figure 4: Fourth-down conversion probability ($y$-axis) according to our model as a function of yards to go ($x$-axis) and the difference in offensive and defensive quality (color).
@@
$\begin{split}{\mathbb{E}}_{\text{go}}[\Delta|{\mathbf{x}},\Delta\geq z]=&\ \vec{\alpha_{1}}\cdot\mathbb{I}\left(\text{fourth down}\right)\cdot\text{B-spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \vec{\alpha_{2}}\cdot\mathbb{I}\left(\text{third down}\right)\cdot\text{B-spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \vec{\alpha_{3}}\cdot\mathbb{I}\left(\text{yards to go}=1\right)\cdot\text{B-spline}\big{(}\text{yardline},\ \text{df}=3\big{)}\\ +&\ \vec{\alpha_{4}}\cdot\mathbb{I}\left(\text{yards to go}\neq 1\right)\cdot\text{B-spline}\big{(}\text{yardline},\ \text{df}=4\big{)}\\ +&\ \beta_{0}+\beta_{1}\cdot\Delta{\mathsf{TQ}}.\end{split}$ (20)
@@
$\begin{split}{\mathbb{E}}_{\text{go}}[\Delta|{\mathbf{x}},\Delta<z]=&\ \alpha_{1}\cdot\mathbb{I}\left(\text{fourth down}\right)\cdot\log(\text{yards to go}+1)\\ +&\ \alpha_{2}\cdot\mathbb{I}\left(\text{third down}\right)\cdot\log(\text{yards to go}+1)\\ +&\ \beta_{0}+\beta_{1}\cdot\Delta{\mathsf{TQ}}.\end{split}$ (21)
@@ 3 Fourth-down decision making under uncertainty
Moreover, the binary win/loss response variable is clustered: all plays from the same game share the same win/loss outcome. This dependence structure reduces the effective sample size, which is somewhere between the number of first-down plays ($229,635$) and the number of non-tied games ($4,101$) in our dataset. According to the simulation study in Brill et al., (2024), we effectively have half as much data as we think. In other words, a win probability estimator fit from a dataset consisting of half as many plays, but where each play has an independent outcome, has the same accuracy as an estimator fit from a clustered dataset like our historical play-by-play dataset.
Thus, there are likely not enough data to experience the full variability of the nonlinear and interacting variables of score differential, time remaining, point spread, yards to opponent endzone, yards to go, timeouts, etc. In fitting win probability models, we are in a limited-data context. Therefore, it is imperative to incorporate uncertainty quantification into the fourth-down decision procedure.
Since we estimate win probability using machine learning models, we need a non-parametric method to quantify uncertainty in these estimates. Bootstrapping is a natural choice to capture such uncertainty. The bootstrapping process begins with generating $B$ bootstrapped datasets from our original dataset. To do so, we use a randomized cluster bootstrap: re-sample games uniformly with replacement and within each game re-sample drives uniformly with replacement. According to the simulation study in Brill et al., (2024), although the randomized cluster bootstrap achieves higher coverage than the standard i.i.d. bootstrap (re-sample plays uniformly with replacement), it produces undercovered intervals that are too narrow. Thus, we also tune our randomized cluster bootstrap using a fractional bootstrap as recommended in Brill et al., (2024). We choose an appropriate value for $B$ ($B=101$) in Appendix C.
@@
Regardless of the effect size, when ${\mathsf{boot\%}}$ is low we cannot trust the point estimate: the effect size at that game-state is too dependent on the random idiosyncrasies of its particular training set. In other words, the edge detected by the point estimate is more due to noise than signal.